# Frustrated Random Walks: A Fast Method to Compute Node Distances on Hypergraphs

### v1

@@ II.2 Frustrated random walks on hypergraphs
In order to measure node distances on hypergraphs, we set one node (call it $t$) as the target, and perform random walks starting from another node (call it $s$). The expected hitting time of this random walk process is interpreted as the distance from $s$ to $t$. Intuitively, the shorter the distance, the closer the relationship between $s$ and $t$. We can think of this random walk process as the donation of a gift. If one person has one gift to donate, she tends to donate it to her close friend. Similarly, when a person is given more than one gift, she also tends to accept her close friend’s. With this observation, we decompose the gift-donating process into two parts: one for proposing the donation, and another for accepting it. By setting an acceptance threshold for the gift donation, we have frustrated the gift donation process, from whence arises our algorithm’s name. Under this scenario, a transition is possible only when a proposal is made and then accepted. Therefore, in frustrated random walks, the transition probability is the product of the proposal probability and the acceptance probability, which is
$T^{F}_{ij,i\neq j}$ $=\frac{\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}{\sum_{j}\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}$ (3)
×$\times\frac{\sum_{\beta}\Big{(}\delta_{\beta}-e_{j\beta}\Big{)}\min\{e_{j\beta},e_{i\beta}\}}{\sum_{k}\sum_{\beta}\Big{(}\delta_{\beta}-e_{j\beta}\Big{)}\min\{e_{j\beta},e_{k\beta}\}}$
@@
Another major difference between simple random walks and frustrated random walks is that in the former case, the proposal to make a transition is always accepted and thus $T^{S}_{ii}=0,\forall i\in\mathbb{V}$, whereas in the latter case, there is a non-zero probability for the proposal to be declined, and thus $T^{F}_{ii}$ is generally non-zero. The diagonal values of $T^{F}$ matrix can be easily calculated from the observation that $\sum_{j}T^{F}_{ij}=1,\forall i,j\in\mathbb{V}$, from which we have
@@
All of these differences combined impact significantly on the calculation of expected hitting times and ultimately the ranking of nodes with respect to a target node. We will show later, using real-world data, that frustrated random walks can better capture human-human interactions in the real world than simple random walks.
II.3 A Unified framework for calculating expected hitting times
In the previous subsections, we outlined simple random walks and frustrated random walks on hypergraphs, and formulated equations to calculate their corresponding transition probabilities. In this section, we extend the calculation to derive the expected hitting times from the transition probabilities. We show that although the transition probabilities of simple and frustrated random walks differ, the calculation of expected hitting times can be formulated in a unified framework.
For a connected hypergraph $H$, denote $N_{t}^{(s)}$ as the hitting time of a random walk process that starts from node $s$, with node $t$ as the target. We use $P(N_{t}^{(s)}=n)$ to denote the probability of reaching target node $t$ from node $s$ after exactly $n$ steps. Since the problem is trivial for the case when $s=t$, we assume the starting node $s$ differs from the target node $t$. Due to the Markov property of the random walk process, whether it be simple or frustrated, we can establish a recurrence equation for the probability, which is
@@ II.3 A Unified framework for calculating expected hitting times
A physical interpretation of the above equation is that we decompose the random walk process into two steps. First, we make a transition from node $s$ to one of its neighbors $i_{s}$; second, due to the Markov property of the random walks, the whole process restarts all over again with $i_{s}$ as the new starting node, with the condition that now the random walker needs to reach the target with exactly $n-1$ steps since one step has already been taken. Thus, Eq. (5) is a recurrence equation for the probabilities $P(N_{t}^{(s)}=n)$. In this equation, we have imposed the condition that $n\geq 2$ because $P(N_{t}^{(j)}=0)=0$ for any node $j\neq t$, and the recursion process terminates immediately when the random walker hits the target. From the transition matrix in Eq. 2 or Eq. 3, we can readily read out the one-step transition probability $P(N_{t}^{(s)}=1)$, which is the initial condition for the recurrence relation in Eq. (5).
Eq. (5) is essentially a system of first order difference equations with constant coefficients. For simplicity of notation, we can understand $N_{t}^{(i)}$ as the $i$th component of a column vector $\boldsymbol{N}_{t}$, and $P(N_{t}^{(i)}=n)$ as the $i$th component of a column vector $\boldsymbol{X}_{n}=P(\boldsymbol{N}_{t}=n)$. With these symbols, we can rewrite Eq. (5) in an abstract form as
@@
Here, $B$ is the coefficient matrix with elements depending on the transition probabilities and the target node $t$. See the appendix A for an illustrative example of how to calculate the $B$ matrix.
II.3.1 Calculation of the $B$ matrix for simple random walks
For simple random walks, $B_{ij}=T^{S}_{ij},\forall i,j\in\mathbb{V}-\{t\}$, and $B_{it}=0,\forall i\in\mathbb{V}-\{t\}$ by definition of Eq. (5). Still from Eq. (5), we have that the left subscript of $B$, which is the index of the starting node for the random walk, can never be equal to $t$. In contrast with the transition matrix $T^{S}$, matrix $B$ is not a Markov matrix because the summation of each row does not always equal one. This non-Markov property of $B$ originates from the existence of the target node $t$. From Eq. (5), we get the rule that if $t$ is not a neighbor of node $i$, then $\sum_{j}B_{ij}=1$, or else $\sum_{j}B_{ij}=1-T^{S}_{it}$. Due to this, the spectral radius of $B$ is always smaller than one.
We have established Eq. (5) to calculate the probability distribution of hitting times using any non-target node $s$ as our starting node. However, in the special case where the starting node $s$ has only one neighbor, and this very neighbor is precisely the target node $t$, then no recurrence equation is needed since we already have $P(N_{t}^{(s)}=n)=\delta_{n,1},\forall n\geq 1$. We call such a node an adherent node of target node $t$, and exclude these nodes from the set of starting nodes (For example, node 4 in Fig. 5 is an adherent to target node 3. ).
@@ II.3.1 Calculation of the $B$ matrix for simple random walks
II.3.2 Calculation of the $B$ matrix for frustrated random walks
The calculation of the $B$ matrix elements for frustrated random walks is similar to that for simple random walks. Here, we still have $B_{ij}=T^{F}_{ij},\forall i,j\in\mathbb{V}-\{t\}$ and $B_{it}=0,\forall i\in\mathbb{V}-\{t\}$. The existence of target node $t$ still renders $B$ a non-Markov matrix in that $\sum_{j}B_{ij}=1$ if $i$ and $t$ do not lie simultaneously in any hyperedge, and $\sum_{j}B_{ij}=1-T^{F}_{it}$ otherwise. The only difference is that due to the imposition of the acceptance threshold for the frustrated random walks, the transition probability from an adherent to the target no longer obeys the sharp $\delta$ distribution, but a geometric distribution, i.e., $P(N_{t}^{(adherent)}=n)=(1-p)^{n-1}p$, where $p$ is the probability of $t$ accepting the transition from the adherent. Consequently, the expected hitting time from the adherent to the target is $1/p$.
II.3.3 Numerical computation of expected hitting times
@@
Here, $\boldsymbol{X}_{1}$ is the probability of hitting the target in the first step, a probability that can be conveniently read out from the transition matrix. By definition, the expected hitting times from an arbitrary node to the target node is
@@
The above equation involves an infinite summation and the powers of the $B$ matrix, and is thus difficult to compute numerically. However, we devise a convenient workaround to convert the above equation into an equivalent linear system of equations which we can efficiently solve using highly optimized linear algebra packages. To do this, we first define the probability generating function, which is
@@
It is straightforward to show that the convergence radius of the above power series is larger than 1. More precisely, we can pinpoint the convergence radius of the series in Eq. (11) at $R=1/\rho(B)$, where $\rho(B)$ is the spectral radius of the coefficient matrix $B$. $\rho(B)$ is guaranteed to be smaller than one by the Gershgorin circle theoremMeyer (2023). With the help of Eq. (8), the probability generating function can be evaluated exactly as
@@
Here, the summation can be evaluated in closed form because the convergence radius of the power series is larger than one. Still due to the fact that convergence radius is larger than 1, we can differentiate the infinite series in Eq. (11) term by term at $z=1$ and obtain an equivalent expression for the expected hitting times as
@@
By definition, $\boldsymbol{f}(z=1)$, which represents the sum of all probabilities, is a column vector where each element is 1. Therefore, to calculate the expected hitting times $\mathbb{E}\boldsymbol{N}_{t}=\boldsymbol{f}^{\prime}(z=1)$, we only need to solve the linear equation (14), which is straightforward to accomplish numerically. It is noteworthy that in Eq. (14), the coefficient matrix $B$ depends both on the hypergraph structure and the target node $t$, and thus target information is implicit in that equation. For real-world hypergraph data, the coefficient matrix in Eq. (14) is generally large and sparse. The best method to numerically solve a large and sparse system is the conjugate gradient methodHestenes and Stiefel (1952) or its variants, which we will use throughout our paper.
Our method for calculating hypergraph node distances can be summarized in Algorithm 1. This algorithm applies to both simple random walks and frustrated random walks. The only difference lies in the calculation of the probability transition matrix. For the sake of simplicity, we only show how to calculate the expected hitting times of simple random walks. The extension of this algorithm to frustrated random walks is straightforward.
@@
3:  Calculate transition matrix $B$ from hypergraph structure and target node $t$, following the procedures in Section II.3.1.
@@
6:  Solve $(I-B)\boldsymbol{x}=\mathbbm{1}$ using conjugate gradient method.
@@
III Experimental results
In the previous sections, we have outlined two random walk algorithms, the simple random walk (SRW) and frustrated random walk (FRW), and showed that we can compute their expected hitting times in a unified framework. In this section, we use real-world hypergraph data sets to show that we can use the expected hitting times as node distances to find each node’s nearest neighbors. We also compare their results with that of DeepWalk.
III.1 Hypergraphs with ground-truth node labels
@@
First, we create hypergraphs using articles published in arXiv under the category of physics from 2017 to 2020. Hypergraph nodes represent articles, and articles written by the same author are collected into the same hyperedge. From the hypergraph, we extract its largest connected component, which contains $52,144$ nodes (articles) and $45,188$ hyperedges (authors). The node and hyperedge degree distributions are shown in Fig. 1. In this data set, the node (i.e., article) degree represents the number of authors of the article. Here, the node degree distribution only approximately follows a power law since in recent literature most papers have more than one author. When we plot the hyperedge (i.e., author) degree distribution, a more visible power law also emerges since there is neither an upper limit nor lower limit on the number of articles that can be authored by the same researcher.
@@
Each arXiv article has one or more subjects, which we can use as its ground truth label. Using the methods described in this paper, we can subsequently calculate hypergraph node distances, and thus find each article’s nearest neighbors. Intuitively, the shorter the distance between two articles, the more similar they should be to each other. We can quantify two articles’ similarity by calculating the Jaccard similarity between their subjects. For comparison, we find each article’s nearest neighbors using simple random walks (SRW), frustrated random walks (FRW) and DeepWalk, and calculate the mean Jaccard similarity between the target article and its top ten nearest neighbors. For one specific article, each method gives a mean Jaccard similarity score. The higher the mean Jaccard similarity, the better the method. We randomly select 12 articles, and for each method, calculate these 12 articles’ average Jaccard similarity score. The average score for each method is shown in Table 1.
@@
From table 1, we can see that FRW has the highest Jaccard similarity (denoted in bold in the table). The arXiv hypergraph is approximately scale-free and lightly weighted. While FRW achieves a better result, We will further show in Section III.2 that the advantage of FRW over SRW is most obvious on heavily weighted scale-free hypergraphs.
The second data set we experiment with is the trivago-clicks data setBenson (2021). The data set has been introduced by the authors of Ref. Chodrow et al. (2021) as a collection of behavior data from users browsing trivago.com while trying to book a hotel. In this hypergraph data set, nodes represent accommodations (mostly hotels) browsed by users, and hyperedges represent a user’s browsing history in the same browsing session before checkout (i.e., before an order is placed). Each node (accommodation) is associated to a country, which can be used as the node’s label. The data set contains $172,738$ nodes and $233,202$ hyperedges. The number of distinct node labels (countries or regions) is $160$. The node degree follows a power law, yet the node degrees are generally small, ranging from $1$ to $339$, i.e., significantly smaller thatn the hypergraph node number. As a result, this hypergraph is still lightly weighted. A log-log plot of the node degree distribution is shown in Fig. 2.
@@
III.2 Heavily-weighted and Scale-free hypergraphs
As already stated, the node degree distribution of a scale-free hypergraph follows a power law. A hypergraph is called heavily weighted if its node degree range spans several orders of magnitude, and the maximum node degree is at least as large as the hypergraph node number. Power law distribution is ubiquitous in real world, e.g., the distributions of wealth among the world population, the number of followers of social media accounts, the citation times of scientific papers, etc. all follow a power law. According to Ref. Barabási and Albert (1999), when constructing a graph following the preferential attachment principle (which refers to attaching a newly created node to an existing node with a probability that is proportional to the node’s degree), the resultant graph will be scale-free. In a scale-free graph or hypergraph, the nodes accumulate their degrees according to the Matthew effect, meaning that the higher a node’s degree, the easier it is for its degree to get even higher. FRW is designed to capture the Matthew effect. Consequently, its advantage is most obvious for heavily-weighted and scale-free hypergraphs, as we will show in this section.
We first apply our results on a hypergraph created from Dream of the Red Chamber, a novel written by Cao Xueqin. We represent novel characters as hypergraph nodes, and each scene in the novel as a hyperedge. Characters are collected into a hyperedge if they appear in the same scene in the novel. Node degree indicates the number of times a node (character) appears in the novel. This hypergraph contains $324$ nodes and $1,524$ hyperedges. The node degree distribution follows a power law, as shown in Fig 3. The node degrees range from $1$ to $954$ which is much larger than the node number, meaning the hypergraph is heavily weighted. We will show that for such a hypergraph, the FRW method is most suitable for calculating node distances.
Unlike the arXiv hypergraph, the novel data set does not have ground truth labels for nodes, which compels us to only compare our random walk results against results obtained from DeepWalk. The fact that DeepWalk results are consistent with human judgement further justifies this results comparison. We first run DeepWalk on the hypergraph to obtain a node ranking with respect to a target node, and then use SRW and FRW methods to get the target’s nearest neighbors. Thus, with respect to each target node, we obtain three rankings using three methods. We then calculate the rank difference between SRW/FRW results and DeepWalk results. To do so, we get the top 10 neighbors of a target node using the SRW/FRW method, find the mean DeepWalk ranking for these 10 neighbors, and finally obtain for each target node the difference between mean DeepWalk rankings and mean SRW/FRW rankings. We summarize the rank differences for several nodes in table 3. To make the results representative enough, we selected three kinds of nodes (characters) from the novel: the central characters of highest node degrees, the supporting characters of medium node degrees, and the marginal characters of low node degrees. It can be seen from the table that for this data set, FRW is much better at finding nearest neighbors of central nodes than SRW. For marginal nodes of low node degrees, the difference between FRW and SRW is negligible.
Target node Jia Baoyu Wang Xifeng Lin Daiyu Jia Zheng Jia Tanchun Jia Huan Jia Yucun Zhen Shiyin Wang Yitie
Node degree 954 554 488 279 188 94 56 22 1
SRW rank diff 99.44 86.33 41.22 74.56 53.33 43.33 0.22 0.0 0.67
FRW rank diff 7.33 0.78 4.11 4.44 4.67 16.44 0.33 0.0 2.33
Table 3: Rank difference between SRW/FRW and DeepWalk for different target nodes in Dream of the Red Chamber data set. Target nodes are ranked from high to low node degrees. For central nodes of high degrees, the differences between FRW and DeepWalk ranking results are much smaller than that for SRW, indicating that the ranking results of FRW are more consistent with DeepWalk results. Yet for marginal nodes of low degrees, the ranking results from SRW, FRW and DeepWalk are equally good.

### v2

@@ III.3 Frustrated random walks on hypergraphs
To address the unique structure of scale-free hypergraphs, we introduce a frustration mechanism to random walks. This approach aims to suppress transitions from peripheral nodes to central nodes while preserving transitions among central nodes. In measuring node distances on hypergraphs using random walks, we designate one node $t$ as the target and initiate random walks from another node $s$. The expected hitting time of this random walk process is interpreted as the distance from $s$ to $t$. Intuitively, the shorter the distance, the closer the relationship between $s$ and $t$. To more accurately measure node distances in real-world hypergraphs, we can conceptualize this random walk process as analogous to the donation of a gift. If one person has one gift to donate, she tends to give it to her close friend. Similarly, when a person is offered multiple gifts, she tends to accept her close friend’s. With this observation, we decompose the gift-donating process into two parts: one for proposing the donation, and another for accepting it. By introducing the acceptance probability to the gift donation, we have frustrated the gift donation process, from whence arises our algorithm’s name. Under this scenario, a transition is possible only when a proposal is made and then accepted. Therefore, in frustrated random walks, the transition probability is the product of the proposal probability and the acceptance probability, which is
$T^{F}_{ij,i\neq j}$ $=\frac{\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{j\alpha}\}}{\sum_{k\neq i}\sum_{\alpha}\Big{(}\delta_{\alpha}-e_{i\alpha}\Big{)}\min\{e_{i\alpha},e_{k\alpha}\}}$ (3)
×$\times\frac{\sum_{\beta}\Big{(}\delta_{\beta}-e_{j\beta}\Big{)}\min\{e_{j\beta},e_{i\beta}\}}{\sum_{l\neq j}\sum_{\beta}\Big{(}\delta_{\beta}-e_{j\beta}\Big{)}\min\{e_{j\beta},e_{l\beta}\}}$
To see the concrete difference between Eq. (2) and Eq. (3), we will use the Harry Potter data set as an example to calculate the one-step transition probabilities from Hokey and Hermione Granger to Harry Potter using these two equations. The results are shown in Table 1. From the table, we can see that in simple random walks, the one-step transition probability from Hokey to Harry Potter is 1.7 times as large as that from Hermione Granger to Harry Potter. The situation reverses in frustrated random walks, for which the one-step transition probability from Hermione Granger to Harry Potter is 289 times larger than that from Hokey to Harry Potter. Since we know that Hermione Granger is a close friend of Harry Potter, the result from frustrated random walks is more consistent with our intuition.
Method name $T_{\text{Hokey}\rightarrow\text{Harry Potter}}$ $T_{\text{Hermione Granger}\rightarrow\text{Harry Potter}}$
Simple random walks $\frac{19}{72}\approx 0.264$ $\frac{9380}{60803}\approx 0.154$
Frustrated random walks $\frac{19}{72}\times\frac{19}{93678}\approx 5.35\times 10^{-5}$ $\frac{9380}{60803}\times\frac{9380}{93678}\approx 0.015$
Table 1: Comparison of one-step transition probabilities from Hokey and Hermione Granger to Harry Potter, in simple and frustrated random walk scenarios. The fact that the transition probability from Hermione Granger to Harry Potter is almost 300 times larger than that from Hokey to Harry Potter in the frustrated random walks authenticates our algorithm, demonstrating its ability to reflect more intuitive relationships.
@@
Another major difference between simple random walks and frustrated random walks is that in the former scenario, the proposal to make a transition is always accepted and thus $T^{S}_{ii}=0,\forall i\in\mathbb{V}$, whereas in the latter scenario, there is a non-zero probability for the proposal to be declined, and thus $T^{F}_{ii}$ is generally non-zero. The diagonal values of $T^{F}$ matrix can be easily calculated from the observation that $\sum_{j}T^{F}_{ij}=1,\forall i,j\in\mathbb{V}$, from which we have
@@
All of these differences combined impact significantly on the computation of expected hitting times and ultimately the ranking of nodes with respect to a target node. We will show later, using real-world data, that frustrated random walks can better capture human-human relationships in the real world than simple random walks.
III.4 A Unified framework for calculating expected hitting times
In the previous sections, we have outlined simple random walks and frustrated random walks on hypergraphs, and formulated equations to calculate their corresponding transition probabilities. In this section, we extend the calculation to derive the expected hitting times from the transition probabilities. We show that although the transition probabilities of simple and frustrated random walks differ, the calculation of expected hitting times can be formulated in a unified framework.
For a connected hypergraph $H$, denote $N_{t}^{(s)}$ as the hitting time of a random walk process that starts from node $s$, with node $t$ as the target. We use $P(N_{t}^{(s)}=n)$ to denote the probability of reaching target node $t$ from node $s$ after exactly $n$ steps. Since the problem is trivial for the case when $s=t$, we demand that the starting node $s$ differ from the target node $t$. Due to the Markov property of the random walk process, whether it be simple or frustrated, we can establish a recurrence equation for the probability, which is
@@ III.4 A Unified framework for calculating expected hitting times
A physical interpretation of the above equation is that we decompose the random walk process into two steps. First, we make a transition from node $s$ to one of its neighbors $i_{s}$; second, due to the Markov property of the random walks, the whole process restarts all over again with $i_{s}$ as the new starting node, on the condition that now the random walker needs to reach the target with exactly $n-1$ steps since one step has already been taken. Thus, Eq. (5) is a recurrence equation for the probabilities $P(N_{t}^{(s)}=n)$. In this equation, we have imposed the condition that $n\geq 2$ because $P(N_{t}^{(j)}=0)=0$ for any node $j\neq t$, and the recursion process terminates immediately when the random walker hits the target. From the transition matrix in Eq. (2) or Eq. (3), we can readily read out the one-step transition probability $P(N_{t}^{(s)}=1)$, which is the initial condition for the recurrence relation in Eq. (5).
Eq. (5) is essentially a system of first order difference equations with constant coefficients. Denoting $N_{t}^{(i)}$ as the $i$th component of a column vector $\boldsymbol{N}_{t}$, and $P(N_{t}^{(i)}=n)$ as the $i$th component of a column vector $\boldsymbol{X}_{n}=P(\boldsymbol{N}_{t}=n)$, we can rewrite Eq. (5) as
@@
Here, $B$ is the coefficient matrix with elements depending on the hypergraph structure, the random walk scenario (simple or frustrated random walk), and the target node $t$. See the appendix A for an illustration of how to calculate the $B$ matrix.
III.4.1 Calculation of the $B$ matrix for simple random walks
For simple random walks with node $t$ as the target, $B_{ij}=T^{S}_{ij}>0,\forall i,j\in\mathbb{V}-\{t\}$, and $B_{it}=0,\forall i\in\mathbb{V}-\{t\}$ by definition of Eq. (5), where $\mathbb{V}$ is the hypergraph vertex set. This means $B$ is a non-negative matrix. Still from Eq. (5), we have that the left subscript of $B$, which is the index of the starting node for the random walk, can never be equal to $t$. In contrast with the transition matrix $T^{S}$, matrix $B$ is not a Markov matrix because the summation of each row does not always equal to one. This non-Markov property of $B$ originates from the existence of the target node $t$. From Eq. (5), we get the rule that
$\sum_{j}B_{ij}=\begin{cases}1&\text{ if }t\text{ is not adjacent to }i\\ 1-T^{S}_{it}<1&\text{otherwise}\end{cases}$ (7)
The connectedness of the hypergraph ensures the irreducibility of matrix $B$. Such a matrix as $B$ is called a substochastic matrix. According to Perron–Frobenius theorem for irreducible and non-negative matrices, the spectral radius of matrix $B$, denoted as $\rho(B)$, satisfies the following inequalityMeyer (2023):
$\min_{i}\sum_{j}B_{ij}\leq\rho(B)\leq\max_{i}\sum_{j}B_{ij}=1$ (8)
It can further be proved that the spectral radius of matrix $B$ must be smaller than one () (https://math.stackexchange.com/users/25136/daniel johnson). A heuristic argument proof of this theorem is given in Appendix B. We will use this fact to quickly compute expected hitting times in this paper.
We have established Eq. (5) to calculate the probability distribution of hitting times using any non-target node $s$ as our starting node. However, in the special case where the starting node $s$ has only one neighbor, and this very neighbor is precisely the target node $t$, then no recurrence equation is needed since we already have $P(N_{t}^{(s)}=n)=\delta_{n,1},\forall n\geq 1$. We call such a node an adherent node of target node $t$, and exclude these nodes from the set of starting nodes in Eq. (5) (For example, node 4 in Fig. 6 is an adherent to target node 3.).
@@ III.4.1 Calculation of the $B$ matrix for simple random walks
III.4.2 Calculation of the $B$ matrix for frustrated random walks
The calculation of the $B$ matrix elements for frustrated random walks is similar to that for simple random walks. Here, we still have $B_{ij}=T^{F}_{ij},\forall i,j\in\mathbb{V}-\{t\}$ and $B_{it}=0,\forall i\in\mathbb{V}-\{t\}$. The existence of target node $t$ still renders $B$ a non-Markov matrix in that $\sum_{j}B_{ij}=1$ if $i$ and $t$ do not lie simultaneously in any hyperedge, and $\sum_{j}B_{ij}=1-T^{F}_{it}<1$ otherwise. This matrix is again a substochastic matrix. However, there are two major differences we want to emphasize. The first is that due to the introduction of the acceptance probability into the frustrated random walks, the transition probability from an adherent to the target no longer obeys the sharp $\delta$ distribution, but a geometric distribution, i.e., $P(N_{t}^{(adherent)}=n)=(1-p)^{n-1}p$, where $p$ is the probability of $t$ accepting the transition from the adherent. Consequently, the expected hitting time from the adherent to the target is $1/p$. The second is that the transition from node $i$ to $j$ may be declined, which means $B_{ii}\neq 0$ for at least one $i\in\mathbb{V}$. This guarantees that $B$ is a primitive matrixMeyer (2023), for which the inequality (8) holds. The same proof in the previous subsection guarantees that $\rho(B)<1$.
III.4.3 Numerical computation of expected hitting times
@@
Here, the vector component $\boldsymbol{X}_{1}^{(i)}$ is the probability of hitting the target starting from node $i$ in the first step, a probability that can be conveniently read out from the transition matrix. By definition, the expected hitting times from an arbitrary node to the target node is
@@
The above equation involves an infinite summation and the powers of the $B$ matrix, and is thus difficult to compute numerically. However, we can devise a convenient method to convert the above summation into an equivalent system of linear equations that we can efficiently solve using highly optimized linear algebra packages. To do this, we first define the probability generating function, which is
@@
It is straightforward to show that the radius of convergence of the above power series is larger than 1. More precisely, we can pinpoint its radius of convergence at $R=1/\rho(B)$, where $\rho(B)$ is the spectral radius of the coefficient matrix $B$. A detailed derivation of the radius of convergence of Eq. (13) can be found in Appendix C. We have already shown that $\rho(B)$ is guaranteed to be smaller than one for either simple or frustrated random walks. With the help of Eq. (10), the probability generating function can be evaluated exactly as
@@
Here, the summation can be evaluated in closed form because the radius of convergence of the power series is larger than one. Still due to the fact that radius of convergence is larger than 1, we can differentiate the infinite series in Eq. (13) term by term at $z=1$ and obtain an equivalent expression for the expected hitting times as
@@
The coefficient matrix in the above equation is invertible due to the fact that $\rho(B)<1$. By definition, $\boldsymbol{f}(z=1)$, which represents the sum of all probabilities, is a column vector each element of which is 1. Therefore, to calculate the expected hitting times $\mathbb{E}\boldsymbol{N}_{t}=\boldsymbol{f}^{\prime}(z=1)$, we only need to solve the linear equation (16), which is straightforward to accomplish numerically. It is noteworthy that in Eq. (16), the coefficient matrix $B$ depends both on the hypergraph structure and the target node $t$, and thus target information is implicit in that equation. For real-world hypergraph data, the coefficient matrix in Eq. (16) is generally large and sparse. Moreover, under the framework of frustrated random walks, the coefficient matrix $I-B$ is positive definite. The best method to numerically solve linear equations with large, sparse and positive definite coefficient matrix is the conjugate gradient methodHestenes and Stiefel (1952). For simple random walks, $I-B$ is no longer positive definite. Fortunately, variants of conjugate gradient method already exist for such non-positive-definite matrices. Given the rapidity and efficacy of conjugate gradient method and its variants in solving large and sparse linear systems, we will use them throughout our paper.
Our method for calculating hypergraph node distances can be summarized in Algorithm 1. This algorithm applies to both simple random walks and frustrated random walks. The only difference lies in the calculation of the probability transition matrix. For sake of simplicity, in Algorithm 1, we only show how to calculate the expected hitting times of simple random walks. The extension of this algorithm to frustrated random walks is straightforward.
@@
3:  Calculate transition matrix $B$ from hypergraph structure and target node $t$, following the procedures in Section III.4.1.
@@
6:  Solve $(I-B)\boldsymbol{x}=\mathbbm{1}$ using a variant of conjugate gradient method.
@@
IV Experimental results
In the previous sections, we have outlined two random walk algorithms, the simple random walk (SRW) and frustrated random walk (FRW), and showed that we can compute their expected hitting times in a unified framework. In this section, we will use real-world hypergraph data sets to show that we can use the expected hitting times as node distances to find each node’s nearest neighbors. We also compare their results with that of DeepWalk.
IV.1 Hypergraphs with ground-truth node labels
@@
First, we create hypergraphs using articles published in arXiv under the category of physics from 2017 to 2020. Hypergraph nodes represent articles, and articles written by the same author are collected into the same hyperedge. From the hypergraph, we extract its largest connected component, which contains $52,144$ nodes (articles) and $45,188$ hyperedges (authors). The node and hyperedge degree distributions are shown in Fig. 2. In this data set, the node (i.e., article) degree represents the number of authors of the article, and the hyperedge (author) degree counts the number of articles written by the author.
@@
Each arXiv article has one or more subjects, which we can use as its ground truth label. Using the methods described in this paper, we can calculate hypergraph node distances, and thus find each article’s nearest neighbors. Intuitively, the shorter the distance between two articles, the more similar they should be to each other. We can quantify two articles’ similarity by calculating the Jaccard similarity between their subjects. For comparison, we find each article’s nearest neighbors using simple random walks (SRW), frustrated random walks (FRW) and DeepWalk, and calculate the mean Jaccard similarity between the target article and its top ten nearest neighbors. For one specific article, each method gives a mean Jaccard similarity score. The higher the mean Jaccard similarity, the better the method. We randomly select 12 articles, and for each method, calculate these 12 articles’ average Jaccard similarity score. The average score for each method is shown in Table 2.
@@
From table 2, we can see that FRW has the highest Jaccard similarity (denoted in bold type in the table). The arXiv hypergraph is approximately scale-free and lightly weighted. While FRW achieves a better result, we will further show in Section IV.2 that the advantage of FRW over SRW is most obvious on heavily weighted scale-free hypergraphs.
The second data set we experiment with is the trivago-clicks data setBenson (2021). This data set has been introduced by the authors of Ref. Chodrow et al. (2021) as a collection of behavior data from users browsing trivago.com while trying to book a hotel. In this hypergraph data set, nodes represent accommodations (mostly hotels) browsed by users, and hyperedges represent a user’s browsing history in the same browsing session before checkout (i.e., before an order is placed). Each node (accommodation) is associated to a country, which can be used as the node’s label. The data set contains $172,738$ nodes and $233,202$ hyperedges. The number of distinct node labels (countries or regions) is $160$. The node degree follows a power law, yet the node degrees are generally small, ranging from $1$ to $339$, which is significantly smaller than the hypergraph node number. As a result, this hypergraph is still lightly weighted. A log-log plot of the node degree distribution is shown in Fig. 3.
@@
In this section, we have shown the effectiveness of FRW in measuring node distances on real-world hypergraphs. Although FRW is effective, its advantage compared to SRW is not yet fully visible. In this next section, we will demonstrate the full advantage of FRW using more heavily-weighted and scale-free hypergraphs.
@@
IV.2 Heavily-weighted and scale-free hypergraphs without node labels
As already stated in the previous sections, the node degree distribution of a scale-free hypergraph that is generated according to the principle of preferential attachment follows the power lawBarabási and Albert (1999). A hypergraph is called heavily weighted if its node degree range spans several orders of magnitude, and the maximum node degree is at least as large as the hypergraph node number. Since we have designed the frustrated random walks (FRW) with preferential attachment in mind, we expect this method to more accurately compute node distances on scale-free hypergraphs, as will be demonstrated in this section.
Unlike the arXiv hypergraph, the hypergraphs in this section do not have ground truth labels for their nodes, which necessitates us to compare our random walk results with that obtained from DeepWalk. We have chosen DeepWalk as our benchmark for three reasons. First, it is easy to generalize DeepWalk from graphs to hypergraphs, whereas for the other node embedding algorithms, such a generalization is highly non-trivial. Second, among various node embedding algorithms, DeepWalk excels at dealing with scale-free networks since it is essentially a language model, and word frequency in natural languages follows the power law distribution (Zipf’s law)Perozzi et al. (2014). Third, the two hypergraphs in this section are both derived from novels written in natural languages, a domain where DeepWalk performs particularly well.